# ACM MMSys 2024 Bandwidth Estimation in Real Time Communications Challenge

### v1

@@ Abstract.
The quality of experience (QoE) delivered by video conferencing systems to end users depends in part on correctly estimating the capacity of the bottleneck link between the sender and the receiver over time. Bandwidth estimation for real-time communications (RTC) remains a significant challenge, primarily due to the continuously evolving heterogeneous network architectures and technologies. From the first bandwidth estimation challenge which was hosted at ACM MMSys 2021, we learnt that bandwidth estimation models trained with reinforcement learning (RL) in simulations to maximize network-based reward functions may not be optimal in reality due to the sim-to-real gap and the difficulty of aligning network-based rewards with user-perceived QoE. This grand challenge aims to advance bandwidth estimation model design by aligning reward maximization with user-perceived QoE optimization using offline RL and a real-world dataset with objective rewards which have high correlations with subjective user-perceived audio/video quality in Microsoft Teams. All models submitted to the grand challenge underwent initial evaluation on our emulation platform. For a comprehensive evaluation under diverse network conditions with temporal fluctuations, top models were further evaluated on our geographically distributed testbed by using each model to conduct calls within a -day period. The winning model is shown to deliver comparable performance to the top behavior policy in the released dataset. By leveraging real-world data and integrating objective audio/video quality scores as rewards, offline RL can therefore facilitate the development of competitive bandwidth estimators for RTC.
@@ 4. Evaluation Setup
The haseline model was trained with a discount factor of , learning rate of , batch size of samples, temparture parameter of , and an expectile parameter of .
@@ 5. Evaluation Results & Discussion
Fast and furious: an actor-critic model which is trained in 2-stages. In the first stage, the critic is independently trained to predict audio and video quality scores. In the second stage, the actor is trained on the emulated dataset in an asymmetric approach which leverages the pre-trained critic and ground truth information in the emulated dataset. Objective audio/video MOS scores are used as rewards.
Schaferct: an actor-critic model which has a GRU cell and residual connections in its architecture. The model has been trained with IQL on about of the testbed dataset. Objective audio/video MOS scores are used as rewards.
@@
SJTU MediaLab: an actor-critic model with a variational auto-encoder (VAE) to learn latent representations of network observation vectors and bandwidth estimates. Model design and training are based on BCQ (Fujimoto et al., 2019) and TD3+BC (Fujimoto and Gu, 2021). A network-metric-based function is used as a reward.
@@
Paramecium: a rule-based bandwidth estimation model that combines delay-based and rate-based congestion control strategies to maintain a small bottleneck queuing delay. No reward function is used for the design.

### v2

@@ Abstract.
The quality of experience (QoE) delivered by video conferencing systems to end users depends in part on correctly estimating the capacity of the bottleneck link between the sender and the receiver over time. Bandwidth estimation for real-time communications (RTC) remains a significant challenge, primarily due to the continuously evolving heterogeneous network architectures and technologies. From the first bandwidth estimation challenge which was hosted at ACM MMSys 2021, we learned that bandwidth estimation models trained with reinforcement learning (RL) in simulations to maximize network-based reward functions may not be optimal in reality due to the sim-to-real gap and the difficulty of aligning network-based rewards with user-perceived QoE. This grand challenge aims to advance bandwidth estimation model design by aligning reward maximization with user-perceived QoE optimization using offline RL and a real-world dataset with objective rewards which have high correlations with subjective audio/video quality in Microsoft Teams. All models submitted to the grand challenge underwent initial evaluation on our emulation platform. For a comprehensive evaluation under diverse network conditions with temporal fluctuations, top models were further evaluated on our geographically distributed testbed by using each model to conduct calls within a -day period. The winning model is shown to deliver comparable performance to the top behavior policy in the released dataset. By leveraging real-world data and integrating objective audio/video quality scores as rewards, offline RL can therefore facilitate the development of competitive bandwidth estimators for RTC.
@@ 4. Evaluation Setup
The baseline model was trained with a discount factor of , learning rate of , batch size of samples, temperature parameter of , and an expectile regression parameter of .
@@ 5. Evaluation Results & Discussion
Fast and furious (Cetinkaya et al., 2024): an actor-critic model which is trained in 2-stages. In the first stage, the critic is independently trained to predict audio and video quality scores. In the second stage, the actor is trained on the emulated dataset in an asymmetric approach which leverages the pre-trained critic and ground truth information in the emulated dataset. Objective audio/video MOS scores are used as rewards.
Schaferct (Tan et al., 2024): an actor-critic model which has a GRU cell and residual connections in its architecture. The model has been trained with IQL on about of the testbed dataset. Objective audio/video MOS scores are used as rewards.
@@
SJTU MediaLab (Lu et al., 2024): an actor-critic model with a variational auto-encoder (VAE) to learn latent representations of network observation vectors and bandwidth estimates. Model design and training are based on BCQ (Fujimoto et al., 2019) and TD3+BC (Fujimoto and Gu, 2021). A network-metric-based function is used as a reward.
@@
Paramecium (Wang et al., 2024): a rule-based bandwidth estimation model that combines delay-based and rate-based congestion control strategies to maintain a small bottleneck queuing delay. No reward function is used for the design.